OpenStudy (anonymous):

I don't know how to do any of the steps for this problem. Can someone please help me? Find the slope-intercept equation of the line with the following properties: Parallel to the line y=−5; containing the point (4,8).

OpenStudy (zzr0ck3r):

the equation will have the form y=mx+b where m is slope and b is the y intercept parallel means it will have the same slope so we have y = -5x+b now we need to solve for b so use the point they gave you (4,8) so when y = 8, x =4 plug this in y=-5x+b becomes 8=-5*4+b solve for b and tell me what you have.

OpenStudy (anonymous):

but y=-5, not m. I thought the same thing and the professor said it wasn't right

OpenStudy (anonymous):

y=8, both are horizontal lines

OpenStudy (anonymous):

but how do you come up with that?

OpenStudy (zzr0ck3r):

ahh its y = -5 not y = -5x sorry so it will be y = 8

OpenStudy (anonymous):

but how do I come to that? I don't understand this and it was due by yesterday.

OpenStudy (zzr0ck3r):

y=-5 has 0 slope so we know the equation will have 0 slope and pass through (4,8) only one line does this y=8

OpenStudy (anonymous):

from the form of the equation y = -5. when y = a constant, it's a horizontal line. no matter what x is, y is always -5 for this line. A horizontal line has a slope of 0. so in y=mx+b, m = 0. thus y = b. in your problem the line parallel to y = -5 goes through the point ( yeah, rendering all this pointless

OpenStudy (anonymous):

Ah. Ok. Thank you very much! :)